OpenStudy (anonymous):

Help please!... Two boys want to use a seesaw, and they need to move the seesaw so that their weights will balance out. The formula is given by w1 • d1 = w2 • d2 where w1 = weight of the first boy, d1 = distance of the first boy from the fulcrum, w2 = weight of the second boy, and d2 = distance of the second boy from the fulcrum. Rewrite the formula to solve for d2. Show all steps in your work.

OpenStudy (anonymous):

we can do that

OpenStudy (anonymous):

\[w_1\times d_1=w_2\times d_2\] \[

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

divide both sides by \(w_2\) and get \[\huge \frac{w_1\times d_1}{w_2}=d_2\]

OpenStudy (anonymous):

divide both sides by \(w_2\) and get \[\huge \frac{w_1\times d_1}{w_2}=d_2\]

OpenStudy (anonymous):

omg im so sorry i did not get back to yall yesterday, my computer was acting up @satellite73 @misty1212